OpenStudy (anonymous):

what are the solutions o the system? y= x^2 + 3x - 4 y= 2x+2

rishavraj (rishavraj):

it means x^2 + 3x - 4 = 2x + 2 ....proceed further

OpenStudy (agent0smith):

Same thing as the last question, set equal to each other, solve for x, then plug in x to find y.

OpenStudy (anonymous):

@agent0smith well i got x=2,-3? what do i plug in to find y?

OpenStudy (agent0smith):

Plug in each of them, one at a time.

OpenStudy (anonymous):

@agent0smith so plug each value into both equations? so basically do like 4 equations? :/

OpenStudy (agent0smith):

No, use one. The easier one. Do the same thing you did in the last question...

rishavraj (rishavraj):

both equations r equal.. so after getting value of x... plug it in any one of the two

OpenStudy (agent0smith):

If you plug each value into both, you'll waste time getting the same y value twice.

OpenStudy (anonymous):

@rishavraj @agent0smith okay well i got x=6?

rishavraj (rishavraj):

no u r wrong

OpenStudy (agent0smith):

"well i got x=2,-3" You already found x. Now you're finding y. Use y= 2x+2. When x=2, y= ? When x=-3, y=?

OpenStudy (anonymous):

WAIT I GOT IT. it's (-3,-4) and (2,6)

OpenStudy (agent0smith):

Yes.